Please help me solve this equation:
x - 1 < = 5 or x + 3 > = 10

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

x - 1 < = 5 or x + 3 > = 10
-----
x <= 6 or x >=7
====================
Cheers,
Stan H.

Answer by mananth(16946)

x - 1 < = 5
add 1
x-1+1<=5+1
x<=6
or x + 3 > = 10
add -3
x+3-3>=10-3
x>=7
...